# Total and Symmetry resolved Entanglement spectra in some Fermionic CFTs from the BCFT approach

### v1

@@ 4.1 Boundary states for $U(1)^{N}$ chiral symmetries
The complex coefficients in eq.(58) are $a_{\lambda}=a_{0}e^{i\theta\cdot\lambda+i\gamma_{R}\left(\lambda\right)}$, where $\theta_{i}$ are the phases similar to the ones appearing in eq.(19)-(20), and the factor $e^{i\gamma_{R}\left(\lambda\right)}$ appears due to a freedom in solutions to the Cardy-Lewellen sewing conditions. The partition function for the boundary conditions preserving the same $U(1)^{N}$ chiral symmetry is given by
@@
where the Affleck-Ludwig g-factor is $g_{R}=a_{0}=\sqrt{\mathrm{Vol}\left(\Lambda[R]\right)}$, the volume of the primitive cell in $\Lambda[R]$, was used to satisfy the cardy conditions. The first equality in eq.(60) gives the partition function in the closed string channel, while the second equality gives the partition function in the open string channel. The lattice $\Lambda^{*}[R]$ is dual of the lattice $\Lambda[R]$. The partition, as written in the open string channel, is diagonal in the $U(1)^{N}$ charge sectors, the charges are given by $\mu+\frac{\Delta\theta}{2}$.
@@
$Z^{\Delta\theta}_{R^{\prime},R}(\tilde{q})=g_{R^{\prime}}g_{R}\tilde{q}^{-\frac{N}{24}}\prod_{m=1}^{\infty}\frac{1}{\det\left(\mathbb{I}-\tilde{q}^{N}R^{\prime T}\cdot R\right)}\sum_{\lambda\in\tilde{\Lambda}[R^{\prime},R]}e^{i\left(\Delta\theta+\pi s\right)\cdot\lambda}\tilde{q}^{\frac{\lambda^{2}}{2}},$ (61)
@@
$\begin{split}Z^{\Delta\theta}_{R^{\prime},R}(\tilde{q})=g_{R^{\prime}}g_{R}\left(\frac{1}{\eta\left(-\frac{1}{\tau}\right)^{M}}\right)\left(\frac{\eta\left(-\frac{2}{\tau}\right)}{\eta\left(-\frac{1}{\tau}\right)}\right)^{N-M-2p}&\left(\frac{1}{\eta\left(-\frac{1}{\tau}\right)^{2p}}\prod_{j=1}^{p}\frac{\vartheta_{1}\left(t_{j}|-\frac{1}{\tau}\right)}{2\sin\left(\pi t_{j}\right)}\right)\\ &\hskip 21.68121pt\sum_{\lambda\in\tilde{\Lambda}[R^{\prime},R]}e^{i\left(\Delta\theta+\pi s\right)\cdot\lambda}\tilde{q}^{\frac{\lambda^{2}}{2}}.\end{split}$ (62)
@@
$\begin{split}Z^{\Delta\theta}_{R^{\prime},R}(\tilde{q})=\frac{g_{R^{\prime}}g_{R}}{\mathrm{Vol}\left(\tilde{\Lambda}[R^{\prime},R]\right)}\sqrt{\mathrm{det}^{\prime}\left(\mathbb{I}-R^{\prime T}\cdot R\right)}&\left(\frac{1}{\eta\left(\tau\right)^{M}}\right)\left(\frac{\eta\left(\tau\right)}{\eta\left(\frac{\tau}{2}\right)}\right)^{N-M-2p}\\ &\left(\prod_{j=1}^{p}\frac{i\eta(\tau)}{\vartheta_{1}\left(t_{j},\tau|\tau\right)}\right)\sum_{Q\in\sigma}q^{\frac{Q^{2}}{2}},\end{split}$ (63)
@@ 4.2 Entanglement Spectra: Total and Symmetry resolved
$\begin{split}S_{1}=N\frac{W}{6}+\log\left(\mathrm{Vol}(\Lambda[R])\right)&-N\sum_{k=1}^{\infty}\left(\log(1-\tilde{q}^{k})-k\log\left(\tilde{q}\right)\frac{\tilde{q}^{k}}{1-\tilde{q}^{k}}\right)\\ &+\log\left(\tilde{q}\right)\frac{\sum_{\lambda\in\Lambda[R]}\lambda^{2}\tilde{q}^{\frac{\lambda^{2}}{2}}e^{i\Delta\theta\cdot\lambda}}{2\sum_{\lambda\in\Lambda[R]}\tilde{q}^{\frac{\lambda^{2}}{2}}e^{i\Delta\theta\cdot\lambda}}+\log\left(\sum_{\lambda\in\Lambda[R]}\tilde{q}^{\frac{\lambda^{2}}{2}}e^{i\Delta\theta\cdot\lambda}\right)\end{split}$ (67)
@@ Appendix A Modular functions
$\vartheta_{1}(z|\tau)$ $=i\sum_{n=-\infty}^{\infty}(-1)^{n}q^{\left(\frac{n^{2}}{2}+\frac{1}{2}\right)}e^{i2\pi n\left(z+\frac{1}{2}\right)}$ (71)
$\vartheta_{2}(z|\tau)$ $=\sum_{n=-\infty}^{\infty}q^{\left(\frac{n^{2}}{2}+\frac{1}{2}\right)}e^{i2\pi n\left(z+\frac{1}{2}\right)}$ (72)
@@
where $q=e^{i2\pi\tau}$, $\tau$ being the modular parameter. The S transformation, that is $\tau\to-\frac{1}{\tau}$, the properties of these functions may be determined using the Poisson re-summation formula, these identities are given below
@@ Appendix B Boundary States for Dirac Fermion
In this appendix, we find the Cardy boundary states for massless Dirac fermions and modular invariant Dirac fermions with either boundary condition given by eq.(19)-(20). The boundary states for the Neveu-Schwartz ($\mathrm{NS}$) sector for either boundary conditions are known, for completeness we give a derivation of these states in both the sectors here.
@@
where $\hat{F}$ is the ladder operator, also called Klein factors, acting on the fermion number eigenstates as $\hat{F}\left|N,\bar{N}\right\rangle=\left|N-1,\bar{N}\right\rangle$. The operator $\hat{N}$ may be thought of as the fermion number operator and $\hat{\lambda}$ is the phase counting operator. The phase counting operator is needed due to the anti-commutation relation of the fermion operators and our convention of the fermion states. The boson operator $e^{i\varphi(z)}$ in eq.(81) conserves the fermion number, but creates excitations on the fixed fermion number ground state. In the Ramond sector, we pick our reference states in the holomorphic sector to be $\left|\sigma\right\rangle$. We also define fermion number operator $\hat{N}$ such that $\hat{N}\left|\sigma\right\rangle=0$ and $\hat{N}\left|\rho\right\rangle=-\left|\rho\right\rangle$ in this sector, it has the relation $\hat{N}=J_{0}-\frac{1}{2}$ with the $U(1)$ current. In the Neveu-Schwartz sector, we have $\hat{N}=J_{0}$ and we have only one choice for the reference ground state. The phase counting operator is defined as
@@ Acknowledgements.
HG is supported by the Prime Minister’s Research Fellowship offered by the Ministry of Education, Govt. of India. HG would also like to thank Urjit A. Yajnik for the helpful discussions, and comments on the manuscript.

### v2

@@ 4.1 Boundary states for $U(1)^{N}$ chiral symmetries
The complex coefficients in eq.(58) are $a_{\lambda}=a_{0}e^{i\theta\cdot\lambda+i\gamma_{R}\left(\lambda\right)}$, where $\theta_{i}$ are the phases similar to the ones appearing in eq.(19)-(20). The factor $e^{i\gamma_{R}\left(\lambda\right)}$ appears in solutions to the Cardy-Lewellen sewing conditions. Since its exact form is not required here, we don’t discuss it further. The partition function for the boundary conditions preserving the same $U(1)^{N}$ chiral symmetry is given by
@@
where the Affleck-Ludwig g-factor is $g_{R}=a_{0}=\sqrt{\mathrm{Vol}\left(\Lambda[R]\right)}$, the volume of the primitive cell in $\Lambda[R]$, was used to satisfy the cardy conditions. The first equality in eq.(60) gives the partition function in the closed string channel, while the second equality gives the partition function in the open string channel. The lattice $\Lambda^{*}[R]$ is dual of the lattice $\Lambda[R]$. The partition, as written in the open string channel, is diagonal in the $U(1)^{N}$ charge sectors and the charges are given by $\mu+\frac{\Delta\theta}{2}$.
@@
$Z^{\Delta\theta}_{R^{\prime},R}=g_{R^{\prime}}g_{R}\tilde{q}^{-\frac{N}{24}}\prod_{m=1}^{\infty}\frac{1}{\det\left(\mathbb{I}-\tilde{q}^{N}R^{\prime T}\cdot R\right)}\sum_{\lambda\in\tilde{\Lambda}[R^{\prime},R]}e^{i\left(\Delta\theta+\pi s\right)\cdot\lambda}\tilde{q}^{\frac{\lambda^{2}}{2}},$ (61)
@@
$\begin{split}Z^{\Delta\theta}_{R^{\prime},R}=g_{R^{\prime}}g_{R}\left(\frac{1}{\eta\left(-\frac{1}{\tau}\right)^{M}}\right)\left(\frac{\eta\left(-\frac{2}{\tau}\right)}{\eta\left(-\frac{1}{\tau}\right)}\right)^{N-M-2p}&\left(\frac{1}{\eta\left(-\frac{1}{\tau}\right)^{2p}}\prod_{j=1}^{p}\frac{\vartheta_{1}\left(t_{j}|-\frac{1}{\tau}\right)}{2\sin\left(\pi t_{j}\right)}\right)\\ &\hskip 21.68121pt\sum_{\lambda\in\tilde{\Lambda}[R^{\prime},R]}e^{i\left(\Delta\theta+\pi s\right)\cdot\lambda}\tilde{q}^{\frac{\lambda^{2}}{2}}.\end{split}$ (62)
@@
$\begin{split}Z^{\Delta\theta}_{R^{\prime},R}=\frac{g_{R^{\prime}}g_{R}}{\mathrm{Vol}\left(\tilde{\Lambda}[R^{\prime},R]\right)}\sqrt{\mathrm{det}^{\prime}\left(\mathbb{I}-R^{\prime T}\cdot R\right)}&\left(\frac{1}{\eta\left(\tau\right)^{M}}\right)\left(\frac{\eta\left(\tau\right)}{\eta\left(\frac{\tau}{2}\right)}\right)^{N-M-2p}\\ &\left(\prod_{j=1}^{p}\frac{i\eta(\tau)}{\vartheta_{1}\left(t_{j},\tau|\tau\right)}\right)\sum_{Q\in\sigma}q^{\frac{Q^{2}}{2}},\end{split}$ (63)
@@ 4.2 Entanglement Spectra: Total and Symmetry resolved
$\begin{split}S_{1}=N\frac{W}{6}+\log\left(\mathrm{Vol}(\Lambda[R])\right)&-N\sum_{k=1}^{\infty}\left(\log(1-\tilde{q}^{k})-k\log\left(\tilde{q}\right)\frac{\tilde{q}^{k}}{1-\tilde{q}^{k}}\right)\\ &+\frac{1}{2}\log\left(\tilde{q}\right)\frac{\sum_{\lambda\in\Lambda[R]}\lambda^{2}\tilde{q}^{\frac{\lambda^{2}}{2}}e^{i\Delta\theta\cdot\lambda}}{\sum_{\lambda\in\Lambda[R]}\tilde{q}^{\frac{\lambda^{2}}{2}}e^{i\Delta\theta\cdot\lambda}}+\log\left(\sum_{\lambda\in\Lambda[R]}\tilde{q}^{\frac{\lambda^{2}}{2}}e^{i\Delta\theta\cdot\lambda}\right)\end{split}$ (67)
@@ Appendix A Modular functions
$\vartheta_{1}(z|\tau)$ $=i\sum_{n=-\infty}^{\infty}(-1)^{n}q^{\frac{1}{2}\left(n-\frac{1}{2}\right)^{2}}e^{i2\pi\left(n-\frac{1}{2}\right)z}$ (71)
$\vartheta_{2}(z|\tau)$ $=\sum_{n=-\infty}^{\infty}q^{\frac{1}{2}\left(n-\frac{1}{2}\right)^{2}}e^{i2\pi\left(n-\frac{1}{2}\right)z}$ (72)
@@
where $q=e^{i2\pi\tau}$, $\tau$ being the modular parameter. The S transformation, that is $\tau\to-\frac{1}{\tau}$, properties of these functions may be determined using the Poisson re-summation formula, these identities are given below
@@ Appendix B Boundary States for Dirac Fermion
In this appendix, we find the Cardy boundary states for massless Dirac fermions and modular invariant Dirac fermions with either boundary condition given by eq.(19)-(20). The boundary states for the Neveu-Schwarz ($\mathrm{NS}$) sector for either boundary conditions are known, for completeness we give a derivation of these states in both the sectors here.
@@
where $\hat{F}$ is the ladder operator, also called Klein factors, acting on the fermion number eigenstates as $\hat{F}\left|N,\bar{N}\right\rangle=\left|N-1,\bar{N}\right\rangle$. The operator $\hat{N}$ may be thought of as the fermion number operator and $\hat{\lambda}$ is the phase counting operator. The phase counting operator is needed due to the anti-commutation relation of the fermion operators and our convention of the fermion states. The boson operator $e^{i\varphi(z)}$ in eq.(81) conserves the fermion number, but creates excitations on the fixed fermion number ground state. In the Ramond sector, we pick our reference states in the holomorphic sector to be $\left|\sigma\right\rangle$. We also define fermion number operator $\hat{N}$ such that $\hat{N}\left|\sigma\right\rangle=0$ and $\hat{N}\left|\rho\right\rangle=-\left|\rho\right\rangle$ in this sector, it has the relation $\hat{N}=J_{0}-\frac{1}{2}$ with the $U(1)$ current. In the Neveu-Schwarz sector, we have $\hat{N}=J_{0}$ and we have only one choice for the reference ground state. The phase counting operator is defined as
@@ Acknowledgements.
HG is supported by the Prime Minister’s Research Fellowship offered by the Ministry of Education, Govt. of India. HG would also like to thank Urjit A. Yajnik for the helpful discussions, and comments on the manuscript. We would also like to thank the anonymous Referee for their insightful comments on the manuscript.